# Rater Cohesion and Quality from a Vicarious Perspective

### v1

@@ 1 Introduction
A crucial part of many AI systems is the humans who provide feedback for learning or evaluation (Vaughan, 2018). As AI systems grow more powerful, aligning models with human values becomes even more critical. Recent work in reinforcement learning with human feedback (RLHF) (Casper et al., 2023; MacGlashan et al., 2023) highlights the gains in model performance from aligning them to human values. This RLHF research also notes the technical challenges associated with doing so.
@@
Are some groups more coherent than others when disclosing their own perceptions of offense?
@@
We address questions using metrics introduced for understanding rater cohesion (Prabhakaran et al., 2023) and CrowdTruth (Dumitrache et al., 2018), two approaches for measuring the impacts of rater disagreement. Rater cohesion metrics measure the extent to which rater disagreement is based on group membership. CrowdTruth teases disagreement apart due to differences of opinion from poor rater quality.
@@ 2 Related Work
Prior work has highlighted the prevalence of disagreement in aggregated labels for subjective NLP tasks such as toxic language detection (Binns et al., 2017; Park et al., 2018; Sap et al., 2019; Davidson et al., 2019; Al Kuwatly et al., 2020). Disagreement is often due to rater identity (race, gender, age, education, and first language) and their beliefs (political leaning) (Sap et al., 2019; Al Kuwatly et al., 2020; Larimore et al., 2021; Sap et al., 2022; Goyal et al., 2022; Pei and Jurgens, 2023; Homan et al., 2023; Weerasooriya et al., 2023; Prabhakaran et al., 2023). Studies have also highlighted the impact of rater bias on NLP datasets (Geva et al., 2019). To uncover and analyze these differences, previous work has relied on regression models and training classifiers using demographic information and comparing their predictions (Binns et al., 2017; Davidson et al., 2019; Al Kuwatly et al., 2020; Larimore et al., 2021; Goyal et al., 2022).
Recent work has advocated the use of non-aggregated (rater-level) labels (Basile et al., 2021; Prabhakaran et al., 2021; Plank, 2022; Cabitza et al., 2023) to enable an extensive treatment of this variation. To this end, Homan et al. (2023) used Bayesian multilevel models to discover intersectional effects between rater demographics and their ratings. Prabhakaran et al. (2023) proposed a framework to analyze (dis)agreement among rater subgroups. CrowdTruth (Dumitrache et al., 2018) is another framework that benefits from rater-level labels for evaluating the quality of a dataset through three dimensions: individual raters, input data items, and the overall dataset.
@@ 3.2 Group Cohesion Metrics
We use the framework and the metrics proposed by Prabhakaran et al. (2023) to compare in-group cohesion and cross-group divergences. This framework utilizes permutation tests along with proposed metrics to measure the variability of judgments by diverse rater subgroups. The metrics are either in-group and cross-group. All metrics are designed so that larger values mean more cohesion.
@@ 3.2.1 In-group Metrics
In-group metrics measure the cohesion among raters within a group. Each metric captures a slightly different aspect of cohesion, and together, they form a robust signal of group cohesion.
$\bullet$   IRR: Inter-rater reliability (IRR) is used to measure agreement among multiple raters in a way that controls for class imbalance in the distribution of ratings over all items. Specifically, we use Krippendorff’s alpha (Krippendorff, 2004), a metric that generalizes several IRR metrics by accepting an arbitrary number of raters, different levels of measurement, handling missing data, and adjusting to small sample sizes.
@@ 3.2.2 Cross-group Metrics
Cross-group metrics measure the cohesion between the raters belonging to different groups. Each of these metrics roughly corresponds to an in-group metric.
@@ 3.3 CrowdTruth
CrowdTruth (CT) (Dumitrache et al., 2018) is a framework that connects the three dimensions: individual raters, input data items, and the dataset. These are interconnected in the CT algorithm to avoid cases where disagreement from low-quality raters can lower the overall data quality as ambiguous or vice-versa. These dimensions connected through the quality of the rater are weighted by the quality of the data items the rater has annotated and the quality of the annotations in the dataset. In this study, we calculate the CT for the entire dataset with the rater demographics and focus on the individual rater quality scores. The relevant score for the research is the worker quality score (WQS), which measures the overall agreement of one rater with other raters. We compute the metrics using a publicly available implementation [^1] https://github.com/CrowdTruth/CrowdTruth-core of CT.
4 Experiments
4.1 Data
We apply this framework to $\mathcal{D}_{\textit{voiced}}$ (Weerasooriya et al., 2023), a collection of 2338 comments on YouTube videos from the official channels of three leading US cable news networks (CNN, FOX, and MSNBC) labeled by diverse raters for offensiveness. Collected over eight years spanning from 2014 to 2022, the comments, therefore, cover a variety of topics. See Tables 89 for more information about the dataset. We consider political leaning and gender as dimensions to compare agreement among different subgroups. This dataset is unique as it also contains labels for vicarious offenses, where raters are asked to predict offenses for others who do not share their political beliefs. We also inspect these vicarious labels of offense and compare the cohesion across subgroups.
In addition, we also apply the framework to the toxicity ratings dataset (Kumar et al., 2021) of 107,620 comments from Twitter, Reddit, and 4chan labeled for toxicity by 17,280 raters. The dataset contains demographic information about the raters and their political leaning but does not contain vicarious annotations. We sample 250 batches of 20 comments, resulting in a total of 5000 items ($\mathcal{D}_{\textit{toxicity}}$ from hereafter). The sampling process ensures that each batch is annotated by the same five raters, consistent with the original dataset. We then remove raters who report multiple political affiliations. See Tables 8 and 10 for more information about the dataset.
@@ 4.2 Examples
We now present three examples from $\mathcal{D}_{\textit{voiced}}$ to highlight our key findings. Section A.3 contains additional examples and further discussion.
@@
The majority vote for EXP1 across groups is offensive. Plurality scores: Democrats: 0.83, Independents: 0.87, Republicans: 0.83, Women: 0.91, Men: 0.75, MenIND: 0.75, WomenIND: 1.0, and MenREP: 0.5. Findings: Even though the comment is targeted towards sons of the President, there is a portion of men as a whole and even republican leaning men finding it not offensive.
@@
The majority vote for EXP2 across groups is offensive. Plurality scores: Democrats: 0.66, Independents: 1, Republicans: 0.78, Women: 0.76, Men: 1.0, and WomenDEM: 0.50. Other genderPP had a score of 1. Findings: In contrast to EXP1, EXP2 contains a targeted attack on the Democrats, yet only 0.66 of the Democrats were cohesive that it is offensive, and women were also less cohesive in contrast to men.
@@
4.3 Results
Tables 1, 2, 3, 5, 6, and 7 show results for in-group and cross-group cohesion for the personal and vicarious offense. Significant results (see Section A.2 for details) are indicated in bold at the $p=0.05$ significance level, $\downarrow$ indicates the result is less than expected under the null hypothesis and $\uparrow$ indicates the result is greater than expected.
4.3.1 Group Cohesion for Personal Offense
Table 1 shows results on $\mathcal{D}_{\textit{voiced}}$ for personal-level offense by political leaning and gender. For each group, we report in-group metrics (indicated by $\cap$’s) for the group and cross-group metrics (indicated by $\otimes$’s) between the group and all raters not in the group. For political leaning, only Independents show uniformly higher in-group and cross-group agreement than other groups. Democrats have significantly lower cross-group agreement but mixed in-group agreement results. They also have the highest GAI score. Republicans have no significant results, but all agreement metrics and GAI are lower than median random scores.
@@ 4.3.1 Group Cohesion for Personal Offense
Table 2 shows results on $\mathcal{D}_{\textit{toxicity}}$ for personal-level offense by political leaning and gender. Republicans show significantly lower IRR, XRR, and voting agreement scores than random groups. Democrats show higher-than-expected in-group scores and lower-than-expected cross-group scores. Independents have higher-than-expected in-group scores. Intersectional groups of political leaning and gender reveal some noteworthy differences. Republican women have higher-than-expected in-group scores with the highest Negentropy (0.664) and Plurality size (0.980) among all groups. In contrast, Republican men have lower-than-expected in- and cross-group scores. Particularly notable is that Republican men have the lowest scores among all groups for all cross-group metrics. They also have the lowest IRR (0.155) among all groups. Democrat men have lower-than-expected cross-group scores with the lowest GAI (0.635). Independent men have higher-than-expected in-group scores with the highest IRR (0.395) and GAI (1.506) among all groups, while Independent women have higher-than-expected cross-group scores with the highest voting agreement (0.356) among all groups.
@@
4.3.2 Group Cohesion for Vicarious Offense
@@ 4.4 CrowdTruth Evaluation
As introduced earlier, CrowdTruth’s triangle of disagreement is dependent on the raters, data item/unit, and dataset/task. We focus on the worker quality score (WQS, ranging from a minimum of 0 to 1) for this study. The WQS measures the overall agreement of one rater over other raters and favors raters that agree with others.
@@ A.3 Metric-Guided Heuristic Study
  EXP1 (MSNBC): Send trump and his deranged softy sons to Iran!!!
The majority vote for EXP1 across groups is offensive. Plurality scores: Democrats: 0.83, Independents: 0.87, Republicans: 0.83, Women: 0.91, Men: 0.75, MenIND: 0.75, WomenIND: 1.0, and MenREP: 0.5.
The majority vote for EXP2 across groups is offensive. Plurality scores: Democrats: 0.66, Independents: 1, Republicans: 0.78, Women: 0.76, Men: 1.0, and WomenDEM: 0.50. Other genderPP had a score of 1.
The majority vote for EXP3 across groups is offensive. Plurality scores: Democrats: 0.66, Independents: 1, Republicans: 0.77, Women: 0.76, Men: 1, WomenDEM: 0.50, and WomenREP: 0.60. Other genderPP had a score of 1.

### v2

@@ 1 Introduction
A crucial part of many AI systems is the humans who provide feedback for learning or evaluation (Vaughan, 2018). As AI systems grow more powerful, aligning models with human values becomes even more critical. Recent work in reinforcement learning with human feedback (RLHF) (MacGlashan et al., 2017; Lin et al., 2020; Ouyang et al., 2022; Casper et al., 2023) highlights the gains in model performance from aligning them to human values. The research into RLHF also notes the technical challenges associated with doing so.
@@
Are some groups more cohesive than others when disclosing their own perceptions of offense?
@@
We address questions using the metrics introduced in the GRASP framework (Prabhakaran et al., 2024) for understanding rater cohesion and CrowdTruth (Dumitrache et al., 2018), two approaches for measuring the impacts of rater disagreement. Rater cohesion metrics measure the extent to which rater disagreement is based on group membership. CrowdTruth teases disagreement apart due to differences of opinion from poor rater quality.
The major takeaways are that, of the political groups, Independents are the most cohesive, both with themselves and with others. Democrats are the least cohesive with others. Republicans are the least internally cohesive.
@@ 2 Related Work
Prior work has highlighted the prevalence of disagreement in aggregated labels for subjective NLP tasks such as toxic language detection (Binns et al., 2017; Park et al., 2018; Sap et al., 2019; Davidson et al., 2019; Al Kuwatly et al., 2020). Disagreement is often due to rater identity (race, gender, age, education, and first language) and their beliefs (political leaning) (Sap et al., 2019; Al Kuwatly et al., 2020; Larimore et al., 2021; Sap et al., 2022; Goyal et al., 2022; Pei and Jurgens, 2023; Weerasooriya et al., 2023; Homan et al., 2024; Prabhakaran et al., 2024). Studies have also highlighted the impact of rater bias on NLP datasets (Geva et al., 2019). To uncover and analyze these differences, previous work has relied on regression models and training classifiers using demographic information and comparing their predictions (Binns et al., 2017; Davidson et al., 2019; Al Kuwatly et al., 2020; Larimore et al., 2021; Goyal et al., 2022).
Recent work has advocated the use of non-aggregated (rater-level) labels (Basile et al., 2021; Prabhakaran et al., 2021; Plank, 2022; Cabitza et al., 2023) to enable an extensive treatment of this variation. To this end, Homan et al. (2024) used Bayesian multilevel models to discover intersectional effects between rater demographics and their ratings. Prabhakaran et al. (2024) proposed GRASP (Group Associations in Perspectives), a framework to analyze (dis)agreement among rater subgroups. CrowdTruth (Dumitrache et al., 2018) is another framework that benefits from rater-level labels for evaluating the quality of a dataset through three dimensions: individual raters, input data items, and the overall dataset.
@@ 3.2 Group Cohesion Metrics
We use the GRASP framework Prabhakaran et al. (2024) and the associated metrics to compare in-group cohesion and cross-group divergences. This framework utilizes permutation tests along with proposed metrics to measure the variability of judgments by diverse rater subgroups. The metrics are either in-group and cross-group. All metrics are designed so that larger values mean more cohesion.
@@ 3.2.1 In-group Metrics
In-group metrics (indicated by $\cap$’s) measure the cohesion among raters within a group. Each metric captures a slightly different aspect of cohesion, and together, they form a robust signal of group cohesion.
$\bullet$   IRR: Inter-rater reliability (IRR) is used to measure agreement among multiple raters in a way that controls for class imbalance in the distribution of ratings over all items. Specifically, we use Krippendorff’s alpha (Krippendorff, 2006), a metric that generalizes several IRR metrics by accepting an arbitrary number of raters, different levels of measurement, handling missing data, and adjusting to small sample sizes.
@@ 3.2.2 Cross-group Metrics
Cross-group metrics (indicated by $\otimes$’s) measure the cohesion between the raters belonging to different groups. Each of these metrics roughly corresponds to an in-group metric.
@@ 3.3 CrowdTruth
CrowdTruth (CT) (Dumitrache et al., 2018) is a framework that connects three key dimensions: individual raters, input data items, and annotations. These dimensions are interconnected in the CT algorithm to prevent situations where disagreement from low-quality raters can reduce the overall data quality or ambiguous data leads to poor rater performance. The quality of the raters is influenced by the quality of the data items they have annotated and the quality of the annotations in the dataset. In this study, we calculate the CT for the entire dataset with the rater demographics and focus on the individual rater quality scores. The relevant score for the research is the worker quality score (WQS), which measures the overall agreement of one rater with other raters. We use a publicly available implementation [^1] https://github.com/CrowdTruth/CrowdTruth-core of CT. [^2] Our code is available at https://github.com/Homan-Lab/rater_cohesion_public/
4 Data
4.1 Datasets
We apply GRASP and CrowdTruth frameworks to $\mathcal{D}_{\textit{voiced}}$ (Weerasooriya et al., 2023), a collection of 2,338 comments on YouTube videos from the official channels of three leading US cable news networks (CNN, FOX, and MSNBC) labeled by diverse raters for offensiveness. Collected over eight years spanning from 2014 to 2022, the comments, therefore, cover a variety of topics R. KhudaBukhsh et al. (2021). Refer to Tables 89 for more information about the dataset. We consider political leaning and gender as dimensions to compare agreement among different subgroups. This dataset is unique as it also contains labels for vicarious offenses, where raters are asked to predict offenses for others who do not share their political beliefs. We also inspect these vicarious labels of offense and compare the cohesion across subgroups.
In addition, we also apply the frameworks to the toxicity ratings dataset (Kumar et al., 2021) of 107,620 comments from Twitter, Reddit, and 4chan labeled for toxicity by 17,280 raters. The dataset contains demographic information about the raters and their political leaning but does not contain vicarious annotations. We sample 250 batches of 20 comments, resulting in a total of 5,000 items ($\mathcal{D}_{\textit{toxicity}}$ from hereafter). The sampling process ensures that each batch is annotated by the same five raters, consistent with the original dataset. We then remove raters who report multiple political affiliations. Refer to Tables 8 and 10 for more information about the dataset.
@@ 4.2 Examples
We now present three examples from $\mathcal{D}_{\textit{voiced}}$ to highlight our key findings. See Appendix A.3 for additional examples and further discussion.
@@
The majority vote for EXP1 across groups is offensive. Plurality scores: Democrats: 0.83, Independents: 0.87, Republicans: 0.83, Women: 0.91, Men: 0.75, MenIND: 0.75, WomenIND: 1.0, and MenREP: 0.5. Findings: Even though the comment is targeted towards sons of the former US President Donald Trump, there is a portion of men as a whole and even republican leaning men finding it not offensive.
@@
The majority vote for EXP2 across groups is offensive. Plurality scores: Democrats: 0.66, Independents: 1, Republicans: 0.78, Women: 0.76, Men: 1.0, and WomenDEM: 0.50. Other genderPP had a score of 1. Findings: In contrast to EXP1, EXP2 contains a targeted attack on the Democrats, yet only 0.66 of the Democrats agreed that it is offensive, and women were less cohesive in contrast to men.
@@
5 Results
Tables 1, 2, 3, 5, 6, and 7 show results for in-group (indicated by $\cap$’s) and cross-group (indicated by $\otimes$’s) cohesion for the personal and vicarious offense. Significant results (see Appendix A.2 for details) are indicated in bold at the $p=0.05$ significance level, $\downarrow$ indicates the result is less than expected under the null hypothesis and $\uparrow$ indicates the result is greater than expected.
5.1 Group Cohesion for Personal Offense
@@
Results on $\mathcal{D}_{\textit{voiced}}$:
Table 1 shows results on $\mathcal{D}_{\textit{voiced}}$ for personal-level offense by political leaning and gender. For each group, we report in-group metrics for the group and cross-group metrics between the group and all raters not in the group. For political leaning, only Independents show uniformly higher in-group and cross-group agreement than other groups. Democrats have significantly lower cross-group agreement but mixed in-group agreement results. They also have the highest GAI score. Republicans have no significant results, but all agreement metrics and GAI are lower than median random scores.
@@
Results on $\mathcal{D}_{\textit{toxicity}}$:
Table 2 shows results on $\mathcal{D}_{\textit{toxicity}}$ for personal-level offense by political leaning and gender. For political leaning, Republicans show significantly lower IRR, XRR, and voting agreement scores than random groups. Democrats show higher-than-expected in-group scores and lower-than-expected cross-group scores. Independents have higher-than-expected in-group scores.
Intersectional groups of political leaning and gender reveal some noteworthy differences. Republican women have higher-than-expected in-group scores with the highest Negentropy (0.664) and Plurality size (0.980) among all groups. In contrast, Republican men have lower-than-expected in- and cross-group scores. Particularly notable is that Republican men have the lowest scores among all groups for all cross-group metrics. They also have the lowest IRR (0.155) among all groups. Democrat men have lower-than-expected cross-group scores with the lowest GAI (0.635). Independent men have higher-than-expected in-group scores with the highest IRR (0.395) and GAI (1.506) among all groups, while Independent women have higher-than-expected cross-group scores with the highest voting agreement (0.356) among all groups.
5.2 Group Cohesion for Vicarious Offense
@@ 5.3 CrowdTruth Evaluation
As introduced earlier, CrowdTruth’s triangle of disagreement is dependent on the raters, data item/unit, and annotations. We focus on the worker quality score (WQS, ranging from a minimum of 0 to 1) for this study. The WQS measures the overall agreement of one rater over other raters and favors raters that agree with others.
@@ 5.3.1 Results of Group Cohesion for Personal Offense after CrowdTruth Filtering
Results on $\mathcal{D}_{\textit{voiced}}$:
Results on $\mathcal{D}_{\textit{voiced}}$:
@@ 8 Limitations
Acknowledgements
We thank the anonymous reviewers for their valuable feedback and suggestions on our work.
Marcos Zampieri is supported by the Virginia Commonwealth Cyber Initiative (CCI) award number N-4Q24-009.